OpenStudy (anonymous):

find the slope of the line. 3x+5y=-2

OpenStudy (triciaal):

rewrite to isolate the y y = mx + b the slope of the line is m

OpenStudy (rulnick):

for Ax + By = C the slope is -B/A

OpenStudy (rulnick):

... your A is 3 and B is 5.

OpenStudy (anonymous):

3x+5y=-2, 3x=-5y-2, 3x+2=-5y, -3/5x-2/5=y I guess you can get it from there

OpenStudy (triciaal):

@rulnick do you mean -A/B

OpenStudy (anonymous):

so the slope is -3/5 right?

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

thanks ♥

OpenStudy (skullpatrol):

You are correct :-) @Megan_xoxo